# Vibe Coding XR: Accelerating AI + XR Prototyping with XR Blocks and Gemini

### 25 Mar 2026 v1

@@ Abstract.
While large language models have accelerated software development through “vibe coding”, prototyping intelligent Extended Reality (XR) experiences remains inaccessible due to the friction of complex game engines and low-level sensor integration. To bridge this gap, we contribute XR Blocks, an open-source, modular WebXR framework that abstracts spatial computing complexities into high-level, human-centered primitives. Building upon this foundation, we present Vibe Coding XR, an end-to-end rapid prototyping workflow that leverages LLMs to translate natural language intent directly into functional XR software. Using a web-based interface, creators can transform high-level prompts (e.g., “create a dandelion that reacts to hand”) into interactive WebXR applications in under a minute. We provide a preliminary technical evaluation on a pilot dataset (VCXR60) alongside diverse application scenarios highlighting mixed-reality realism, multi-modal interaction, and generative AI integrations. By democratizing spatial software creation, this work empowers practitioners to bypass low-level hurdles and rapidly move from “idea to reality.” Code and live demos are available at https://github.com/google/xrblocks and http://xrblocks.github.io/gem.
@@ 1. Introduction
Recent advances in Large Language Models (LLMs) (Vaswani et al., 2017; Brown et al., 2020; Team et al., 2023; Guo et al., 2025) and agentic workflows (Google, 2025b, 2026; Cursor, 5) are fundamentally reshaping software engineering and creative computing. We are witnessing the emergence and growing prevalence of “vibe coding” (Edwards, 2025), a paradigm in which high-level human intent is translated directly into functional software. While tools such as Gemini Canvas (Google, 2025), Antigravity (Google, 2026), Cursor (Cursor, 5), and Claude Code (Anthropic, 2026) have successfully democratized this capability for 2D and even 3D web development (Vichare et al., 2025; Earle et al., 2025), the domain of Extended Reality (XR) remains largely inaccessible. Prototyping intelligent spatial experiences usually requires navigating a fragmented ecosystem of perception pipelines (Zhu et al., 2025; Shi et al., 2025; He et al., 2023), low-level sensor API integrations (Li et al., 2025b; Fang et al., 2020), complex game engines (Unity, 2022; Unreal, 2022), and on-device testing. Crucially, the absence of “vibe coding” in XR stems from a fundamental lack of high-level building blocks at the disposal of LLMs. Current models struggle to generate reliable spatial software because they are forced to reason over low-level, disjointed API plumbing rather than semantic, human-centric spatial concepts.
@@
To bridge this gap, we introduce XR Blocks (Li et al., 2025a), an open-source WebXR framework designed specifically to make spatial computing intelligible to both human creators and generative AI. As the foundational contribution of this work, XR Blocks provides these missing high-level semantic abstractions. It introduces a modular “Reality Model” that treats users, the physical environment, and intelligent agents as first-class, configurable primitives. By absorbing the incidental complexity of sensor fusion and pipeline integration, XR Blocks equips LLMs with a concise, robust vocabulary to generate complex spatial behaviors without hallucinating low-level implementation details.
@@
Building directly upon this enabling framework, we present Vibe Coding XR as our second key contribution: an end-to-end rapid prototyping workflow that leverages the long-context reasoning capabilities of Gemini to act as an expert XR designer. Using a specialized system prompt, XR Blocks templates, and a web-based interface (XR Blocks Gem), this workflow translates natural language prompts directly into functional, physics-aware XR applications in under a minute. Creators can rapidly iterate on these generated scripts within a desktop “simulated reality” environment, or prompt directly within an Android XR (Google, 2025a) headset to test their creations with live hand interaction and environmental sensing.
In summary, we contribute:
@@
The XR Blocks framework, an open-source, modular WebXR architecture that provides the missing high-level building blocks, abstracting low-level perception and interaction pipelines into semantic primitives optimized for AI + XR prototyping.
@@
Vibe Coding XR, a web-based, open-access workflow that leverages LLMs and the XR Blocks framework to translate natural language intent into deployable, physics-aware XR applications.
Evaluation and scenarios of Vibe Coding XR using our VCXR60 dataset, alongside diverse application scenarios demonstrating immersive education, physics-aware interaction, and rapid game prototyping.
@@ 2. Related Work
2.1. GenXR: Creating AI + XR Experiences
The democratization of XR development has evolved from early toolkits like VR Juggler (Bierbaum et al., 2001) and ARToolKit (Kato and Billinghurst, 1999) to comprehensive game engines such as Unity (Unity, 2022) and Unreal (Unreal, 2022). While these native engines offer high ceilings for fidelity, they impose steep learning curves, significant friction for rapid prototyping for XR, and high barriers for sharing executable code. Conversely, WebGL libraries like three.js (three.js authors, 2022) and frameworks like A-Frame (A-Frame Authors, 2025) provide cross-platform accessibility but lack the high-level abstractions necessary for complex XR interaction and seamless AI integration. XR Blocks bridges this divide. By offering a high-level, web-optimized interaction model akin to MRTK (Authors, 2025), VRTK (VRTK Authors, 2025), XRI (Unity, 2025) for Unity, our framework abstracts low-level sensor and perception pipelines so creators can focus on the what of an experience rather than the how. Crucially, as a web-native architecture, XR Blocks makes AI + XR prototyping inherently accessible and reproducible, empowering the community to seamlessly share, fork, and build upon each other’s creations. Vibe Coding XR further leverages this framework to create behavioral AI + XR experiences, with photorealistic 3D content creation (Poole et al., 2022; Tang et al., 2024; Xu et al., 2024; Hu et al., 2025b) remaining out of scope.
@@ 2.2. Vibe Coding: Gaps from AI to XR
The AI community benefits from a “flywheel effect” driven by open ecosystems such as Hugging Face (Hugging Face, 2025) and TensorFlow Hub (Google, 2025), benchmarks like LMArena (Chiang et al., 2024), and composable frameworks including JAX (Bradbury et al., 2018), PyTorch (Paszke et al., 2019), and TensorFlow (ten, 2015). XR, however, lacks a comparable substrate for rapid, intent-driven iteration. Recent efforts have begun to address this by leveraging LLMs to generate spatial objects and scenes in Unity, such as LLMR (De La Torre et al., 2024), DreamCodeVR (Giunchi et al., 4), and Thing2Reality (Hu et al., 2025b). Furthermore, recent systems like DreamGarden (Earle et al., 2025) have demonstrated using LLM-driven hierarchical planning to generate functional 3D game environments, assets, and C++ logic in Unreal Engine.
@@
While these approaches successfully push the boundaries of automated game design and 3D scene composition, they often rely on asynchronous compilation workflows within heavy, closed-engine ecosystems. This hinders the collective “vibe coding” paradigm (Edwards, 2025), in which an agent’s utility is fundamentally tied to its ability to synthesize and share knowledge across the entire relevant domain. To this end, we are motivated to develop a spatial computing platform on the open web inspired by (Du et al., 2023). By wrapping LLMs (Zhou et al., 2025) around the XR Blocks framework, Vibe Coding XR empowers creators to generate immersive, intelligent, and physics-aware interactions in real-time. This web-based approach ensures that every contribution is (1) built upon a transparent, web-standard codebase, (2) designed for rapid iteration, and (3) instantly distributable across diverse devices and users. By prioritizing these pillars, we transition toward an ecosystem where the community does not merely consume agents but actively drives their evolution, fueling a true “flywheel effect” for XR development.
3. Vibe Coding XR
@@
We introduce Vibe Coding XR, a rapid prototyping workflow that combines the generative reasoning capabilities of LLMs with a high-level modular XR framework. This synergy allows creators to bypass the friction of low-level systems programming and focus purely on sculpting spatial behaviors, AI logic, and interactions.
3.1. XR Blocks Framework
At the foundation of our system is XR Blocks, an open-source modular SDK designed to accelerate human-centered AI + XR innovation. Built upon accessible web technologies like WebXR (authors, 2022), three.js (three.js authors, 2022), and TensorFlow.js (ten, 2015), the framework introduces a high-level Reality Model. This unified abstraction treats the user, the physical world, and intelligent agents as first-class, configurable primitives rather than disjoint data streams. The following snippets demonstrate XR Blocks’s design principle to keep things simple: it creates an object at user’s eye level with preferred distance to object, which changes color when user performs a hand pinch or desktop click.
The framework’s architecture was iteratively refined through the development of 34 open-source templates (examples: Figure 4), incorporating feedback from weekly workshops and technical demonstrations.
3.2. Vibe Coding XR Workflow
Vibe Coding XR leverages the long-context capabilities and advanced thinking processes of modern LLMs like Gemini (Team et al., 2023) to function as an expert XR designer and engineer. We developed a specialized system prompt in Appendix C that “teaches” the LLM the XR Blocks architecture through an open-source web interface111XR Blocks Gem: http://xrblocks.github.io/gem; XR Blocks Prompt: http://xrblocks.github.io/prompts. This context includes:
Persona & Guidelines: Establishes the LLM as a domain expert adhering to best practices for room-scale XR environments (e.g., spatial layout, human-scale proportions, and interaction distances).
Package Management: Specifies how dependencies within XR Blocks are handled to enforce recommended default styles.
Source Code & Templates: Grounds the model with the full source code of the core xrblocks.js library, alongside curated templates and samples (illustrated in Figure 4). This strict grounding minimizes API hallucination and encourages adherence to established design patterns.
Users interact with Vibe Coding XR through high-level natural language prompts (or “vibes”). The LLM translates this intent directly into functional XR Blocks scripts (Zhou et al., 2025), which can be previewed immediately in a desktop simulated reality environment or deployed directly to an Android XR headset.
@@ 3.3. Application Scenarios: From Prompt to XR
To demonstrate the versatility of the Vibe Coding XR workflow, we present several prototypes generated via Vibe Coding XR during four 1-hour workshops with 20 participants. These examples (with prompts provided in Appendix B and the supplementary video) highlight the system’s ability to seamlessly integrate the core pillars of the XR Blocks framework: AI + XR capabilities, XR Interaction, and XR Realism.
@@
xx.
@@
As shown in Figure 5, Vibe Coding XR excels at rapidly prototyping interactive learning environments. Prompted with “visualize Euler’s theorem in geometry and explain vertices, edges, and facets” (AS1), it generated an immersive Math Tutor application. The LLM autonomously selected geometries (a tetrahedron, cube, and octahedron) and allowed users to pinch them to trigger different highlighting strategies. Similarly, prompting for an Immersive Chemistry (AS3) resulted in an app featuring educational cards and 3D volumetric visual effects, allowing users to safely observe simulated reactions (e.g., igniting methane or ethylene) via hand-pinch gestures.
@@
By grounding the LLM in XR Blocks’s physics and depth modules, users can generate highly tactile experiences. Prompting for an XR Sports (AS4) application (“Let me play volleyball with hands and collide with my environment”) yielded a textured ball that reacts realistically to both the user’s physical hands and the surrounding room geometry. Furthermore, a Physics Lab prompt successfully implemented a mixed-reality scale where users pick and drop labeled weights to intuitively learn mechanics.
Rapid Game Prototyping & Procedural Generation
@@
This open-source workflow significantly reduces the time required to prototype XR games. Prompted to “create the Chrome Dino game in XR,” the system generated a voxelized version of the classic game—complete with rushing cacti on a semi-transparent lane and audio cues—reducing development time from hours to minutes. Through iterative refinement, creators have also used the system for advanced procedural generation, including rendering immersive NASA starmaps, procedural city layouts, and explorable ancient Egyptian pyramids directly in the headset.
4. Preliminary Technical Evaluation

### 03 Apr 2026 v2

@@ Abstract.
While large language models (LLMs) have accelerated 2D software development through intent-driven “vibe coding”, prototyping intelligent Extended Reality (XR) experiences remains a major challenge. The fundamental barrier is not just the steep learning curve for human creators, but that low-level sensor APIs and complex game engine hierarchies are ill-suited for LLM reasoning, routinely exceeding context windows and inducing syntax hallucinations. To bridge this gap, we contribute XR Blocks, an open-source, LLM-native WebXR framework. Unlike traditional engines, XR Blocks introduces a semantic “Reality Model” that aligns spatial computing primitives (users, physical environments, and agents) with natural language, providing a robust, concise vocabulary optimized for generative AI. Building upon this foundation, we present Vibe Coding XR, an end-to-end prototyping workflow that leverages LLMs to translate high-level prompts (e.g., “create a dandelion that reacts to my hand”) directly into functional, physics-aware mixed-reality applications. To minimize the friction of on-device testing, the workflow introduces a seamless desktop “simulated reality” to headset deployment loop. Finally, we introduce VCXR60, a pilot dataset of 60 XR prompts paired with an automated evaluation pipeline. Our technical evaluation demonstrates high one-shot execution success, enabling practitioners to bypass low-level hurdles and rapidly move from “idea to reality.” Code and live demos are available at https://github.com/google/xrblocks and http://xrblocks.github.io/gem.
@@ 1. Introduction
Recent advances in Large Language Models (LLMs) (Vaswani et al., 2017; Brown et al., 2020; Team et al., 2023; Guo et al., 2025) and agentic workflows (Google, 2025b, 2026; Cursor, 5) are fundamentally reshaping software engineering and creative computing. We are witnessing the growing prevalence of “vibe coding” (Edwards, 2025), a paradigm in which high-level human intent is translated directly into functional software. While tools such as Gemini Canvas (Google, 2025), Antigravity (Google, 2026), Cursor (Cursor, 5), and Claude Code (Anthropic, 2026) have successfully expanded this capability for 2D and, more recently, 3D web development (Vichare et al., 2025; Earle et al., 2025), the domain of Extended Reality (XR) remains largely inaccessible. Prototyping intelligent spatial experiences still requires navigating a fragmented ecosystem of perception pipelines (Zhu et al., 2025; Shi et al., 2025; He et al., 2023), low-level sensor APIs (Li et al., 2025b; Fang et al., 2020), massive game engine hierarchies (Unity, 2022; Unreal, 2022), and the physical friction of constant on-device testing.
Crucially, the absence of “vibe coding” in XR stems from a fundamental mismatch between how spatial software is currently built and how LLMs reason. Current models struggle to generate reliable XR applications because they are forced to operate over massive, low-level syntax trees and fragmented API layers rather than semantic, human-centered spatial concepts. This incidental complexity routinely exhausts context windows, fragments reasoning steps, and induces code hallucinations. To enable reliable intent-driven spatial computing, we need an abstraction layer explicitly designed not just for human accessibility, but for machine reasoning.
@@
To bridge this gap, we introduce XR Blocks 111XR Blocks: https://xrblocks.github.io (Li et al., 2025a), an open-source WebXR (authors, 2022) framework designed specifically to make spatial computing intelligible to generative AI. As the foundational contribution of this work, XR Blocks introduces an LLM-native “Reality Model” that treats the user, the physical environment, and intelligent agents as first-class, configurable primitives. By encapsulating the complexities of sensor fusion, rendering, and spatial physics into a concise, semantic vocabulary, XR Blocks provides LLMs with the precise architectural constraints necessary to generate complex spatial behaviors without hallucinating underlying implementation.
@@
Building upon this enabling framework, we present Vibe Coding XR, an end-to-end rapid prototyping workflow that leverages the long-context reasoning of LLMs to act as an expert XR designer. Using a specialized system prompt and a web-based interface222XR Blocks Gem: https://xrblocks.github.io/gem, this workflow translates natural language directly into deployable, physics-aware XR applications in under 90 seconds. To solve the persistent HCI challenge of spatial testing friction, Vibe Coding XR allows creators to rapidly iterate on generated scripts within a desktop “simulated reality” environment before deploying instantly to an Android XR (Google, 2025a) headset for live hand interaction and environmental sensing.
Finally, evaluating the viability of generated XR software requires moving beyond subjective manual testing. We contribute the VCXR60 dataset, a pilot dataset consisting of 60 diverse spatial computing prompts sourced from four workshops. Coupled with an automated, headless-browser evaluation pipeline, we provide a quantitative baseline for GenXR code generation. Our evaluation, alongside diverse application scenarios demonstrating XR realism and multi-modal interaction, illustrates the robustness of the system. In summary, we contribute:
@@
The XR Blocks Framework: An open-source, LLM-native WebXR architecture that abstracts low-level spatial perception and interaction pipelines into a semantic “Reality Model” optimized for reliable AI code generation.
@@
Vibe Coding XR Workflow: A web-based, rapid prototyping loop that leverages LLMs to translate natural language intent into deployable XR applications, utilizing a novel desktop-simulator-to-headset testing pipeline to reduce iterative friction.
The VCXR60 Dataset & Evaluation: A preliminary dataset and automated testing pipeline for evaluating intent-driven XR code generation, demonstrating the high one-shot viability of our framework across varied applications.
@@ 2. Related Work
2.1. GenXR: Creating AI + XR Experiences
XR development tooling has evolved from early toolkits like VR Juggler (Bierbaum et al., 2001) and ARToolKit (Kato and Billinghurst, 1999) to comprehensive game engines such as Unity (Unity, 2022) and Unreal (Unreal, 2022). While these native engines offer high ceilings for fidelity, they impose steep learning curves, significant friction for rapid prototyping for XR, and high barriers for sharing executable code. Conversely, WebGL libraries like three.js (three.js authors, 2022) and A-Frame (A-Frame Authors, 2025) provide cross-platform accessibility, but lack the high-level abstractions necessary for complex XR interaction and seamless AI integration. XR Blocks bridges this divide. By offering a high-level, web-optimized interaction model akin to MRTK (Authors, 2025), VRTK (VRTK Authors, 2025), XRI (Unity, 2025) for Unity, our framework abstracts low-level sensor and perception pipelines so creators can focus on the what of an experience rather than the how. As a web-native architecture, XR Blocks makes AI + XR prototyping inherently accessible and reproducible, enabling the community to share, fork, and build upon each other’s creations. Vibe Coding XR further leverages this framework to create behavioral AI + XR experiences, with 3D assets creation (Poole et al., 2022; Tang et al., 2024; Xu et al., 2024; Hu et al., 2025b) remaining out of scope.
@@ 2.2. Vibe Coding: Gaps from AI to XR
The AI community benefits from a “flywheel effect” driven by open ecosystems such as Hugging Face (Hugging Face, 2025) and TensorFlow Hub (Google, 2025), benchmarks like LMArena (Chiang et al., 2024), and composable frameworks including JAX (Bradbury et al., 2018), PyTorch (Paszke et al., 2019), and TensorFlow (ten, 2015). XR, however, lacks a comparable substrate for rapid, intent-driven iteration. Recent efforts have begun to address this by leveraging LLMs to generate spatial objects and scenes in Unity, such as LLMR (De La Torre et al., 2024), DreamCodeVR (Giunchi et al., 4), and Thing2Reality (Hu et al., 2025b), with commercial tools such as Bezi (Bez, [n. d.]) pursuing a similar direction. On the infrastructure side, Ubiq-Genie (Numan et al., 2023) proposed a server-client architecture that exposes AI services to Unity-based XR applications. Furthermore, recent systems like DreamGarden (Earle et al., 2025) have demonstrated the use of LLM-driven hierarchical planning to generate functional 3D game environments, assets, and C++ logic in Unreal Engine.
@@
While these approaches successfully push the boundaries of automated game design and 3D scene composition, they often rely on asynchronous compilation workflows within heavy, closed-engine ecosystems. This limits the collective “vibe coding” paradigm (Edwards, 2025), in which an agent’s utility is fundamentally tied to its ability to synthesize and share knowledge across the entire relevant domain. To this end, we develop a spatial computing platform on the open web inspired by (Du et al., 2023). By coupling LLMs (Zhou et al., 2025) directly with the XR Blocks Reality Model, Vibe Coding XR enables creators to generate immersive, intelligent, and physics-aware interactions. This web-based approach ensures that every contribution is (1) built upon a transparent, web-standard codebase, (2) designed for rapid iteration, and (3) instantly distributable across diverse devices and users. By prioritizing these pillars, we move toward an ecosystem where the community does not merely consume agents but actively drives their evolution, supporting a “flywheel effect” for XR.
3. System Architecture
@@
We introduce Vibe Coding XR, a rapid prototyping workflow that combines the generative reasoning capabilities of LLMs with a high-level modular XR framework. This combination allows creators to bypass the friction of low-level systems programming and focus on designing spatial behaviors, AI logic, and interactions.
3.1. The XR Blocks Reality Model
At the foundation of our system is XR Blocks, an open-source LLM-native abstraction layer designed to make spatial computing intelligible to generative AI. Traditional game engines and low-level WebGL APIs represent scenes as massive, deeply nested syntax trees and fragmented data streams. When forced to navigate these structures, LLMs routinely exhaust their context windows, lose track of spatial relationships, and hallucinate API calls. To solve this, XR Blocks introduces a semantic Reality Model. This unified abstraction aligns spatial computing primitives (the user, the physical environment, and intelligent agents) directly with natural language concepts, providing a robust, concise vocabulary optimized for machine reasoning. The following snippet demonstrates XR Blocks’s design principle of keeping things simple and semantically dense for the LLM: it creates an object at the user’s eye level with a preferred distance to the object, which changes color when the user performs a hand pinch or desktop click.
@@ 3.1. The XR Blocks Reality Model
The framework’s architecture was iteratively refined through the development of 34 open-source templates (examples: Figure 4), incorporating feedback from weekly workshops and live demos.
3.2. System Prompt Architecture
Vibe Coding XR leverages the long-context capabilities and advanced thinking processes of modern LLMs like Gemini (Team et al., 2023) to function as an expert XR designer and engineer. Merely providing an API is insufficient; the LLM must be constrained and guided. We developed a specialized system prompt architecture in Appendix C that explicitly “teaches” the LLM the XR Blocks Reality Model through an open-source web interface. This includes:
Persona & Guidelines: Configures the LLM with a domain expert persona adhering to best practices for room-scale XR environments (e.g., spatial layout, human-scale proportions, and interaction distances).
Package Management: Specifies how dependencies within XR Blocks are handled to enforce recommended styles.
Source Code & Templates: Grounds the model with the full source code of the core xrblocks.js library, alongside curated templates and samples (illustrated in Figure 4). This strict grounding acts as a technical constraint, minimizing API hallucination and enforcing adherence to established spatial design patterns.
3.3. The Simulated-to-Extended Reality Loop
Users interact with Vibe Coding XR through high-level natural language prompts (or “vibes”). The LLM translates this intent directly into functional XR Blocks scripts (Zhou et al., 2025). Crucially, to solve the persistent HCI friction of iterative spatial testing, where creators must constantly don and doff a headset, the workflow introduces a seamless simulated to extended reality loop. Scripts can be previewed immediately in a desktop simulated reality environment, allowing for rapid visual validation of physics, logic, and layout before being deployed directly to an Android XR headset for real-world interaction testing.
@@ 3.4. Application Scenarios: From Prompt to XR
To demonstrate the versatility of the Vibe Coding XR workflow, we present several prototypes generated via Vibe Coding XR:
@@
The image is a grid of sixteen screenshots divided into four horizontal rows. Each row consists of a four-part sequence showing a first-person perspective of a user interacting with digital elements in a virtual or augmented space. The top row, a user’s hands manipulating various 3D geometric shapes—a pyramid, a wireframe tetrahedron, an octahedron, and a cube—while a blackboard in the background displays mathematical formulas. The second row, a user interacts with a balance scale and colorful weights to explore physical principles. The third row, a digital lesson on the ”Combustion of Hydrocarbons,” featuring virtual Bunsen burners with animated orange and blue flames beneath a large informational display. The final row, a person in a living room environment attempting to hit or track a small, multicolored ball as it flies through the air, with red circular reticles appearing to guide the user’s aim or focus.”
@@
As shown in Figure 5, Vibe Coding XR supports rapid prototyping of interactive learning environments. Prompted with “visualize Euler’s theorem in geometry and explain vertices, edges, and facets” (AS1), it generated an immersive Math Tutor application. The LLM autonomously selected geometries (a tetrahedron, cube, and octahedron) and allowed users to pinch them to trigger different highlighting strategies. Similarly, prompting for an Immersive Chemistry (AS3) application produced an experience featuring educational cards and 3D volumetric visual effects, allowing users to safely observe simulated reactions (e.g., igniting methane or ethylene) via hand-pinch gestures.
@@
By grounding the LLM in XR Blocks’s physics and depth modules, users can generate highly tactile experiences. Prompting for an XR Sports (AS4) application (“Let me play volleyball with hands and collide with my environment”) yielded a textured ball that reacts realistically to both the user’s physical hands and the surrounding room geometry. Furthermore, a Physics Lab prompt (AS2) successfully implemented a mixed-reality scale where users pick and drop labeled weights to intuitively learn mechanics.
Game Prototyping & Procedural Generation
@@
This workflow substantially reduces the time required to prototype XR games. Prompted to “create the Chrome Dino game in XR,” (AS5) the system generated a voxelized version of the classic game, complete with rushing cacti on a semi-transparent lane and audio cues, reducing development time from hours to minutes.
4. Preliminary Evaluation with VCXR60